OpenStudy (anonymous):

There are 20 litres of a 65% alcoholic drink. How many litres of a 77% solution of alcohol must be added to obtain a 68% alcoholic drink?

OpenStudy (aravindg):

Okay so first find amount of alcohol in 20 litres. Can you do that?

OpenStudy (anonymous):

13 liters alochol and 7 litres

OpenStudy (anonymous):

why u deleted that

OpenStudy (anonymous):

@douglaswinslowcooper

OpenStudy (aravindg):

Now next we can assume we are adding 100 litre of solution containing alcohol 77 litre alohol.

OpenStudy (anonymous):

I gave too much of the answer and was warned not to do that, so it was deleted.

OpenStudy (aravindg):

This means the net alcohol amount is now 13+77= Total amount of alcohol 7+23=part of solution not containing alcohol

OpenStudy (anonymous):

100,30

OpenStudy (aravindg):

Or wait...It would be better we bring variable initially itself. 13+(77/100)x=Net alcohol content Now this alcohol content should account for 68% 13+(77/100)x ------------ x 100=68 20+x

OpenStudy (aravindg):

Solve for x

OpenStudy (anonymous):

x=6.7

OpenStudy (aravindg):

Good work!!

OpenStudy (anonymous):

i have another question

OpenStudy (aravindg):

Sure post it separately. The other reply was deleted as we are not permitted to post direct answers in this site like yahoo answers. For more info see: https://openstudy.com/code-of-conduct